# On automorphism groups of smooth hypersurfaces

### v1

@@ 1. Introduction
The purpose of this paper is to study smooth hypersurfaces in complex projective spaces with large automorphism groups. Let $F$ in ${\mathbb{C}}[x_{1},x_{2},\ldots,x_{n+2}]$ be a smooth homogeneous polynomial of degree $d\geq 3$ and let $X_{F}\subset{\mathbb{P}}^{n+1}$ denote the smooth hypersurface defined by $F$, where $n\geq 1$. An automorphism $f\in{\rm Aut}(X_{F})$ of $X_{F}$ is called linear with respect to the embedding $X_{F}\hookrightarrow{\mathbb{P}}^{n+1}$ if $f$ extends to an automorphism of ${\mathbb{P}}^{n+1}$, i.e., $f$ is given by a linear change of homogeneous coordinates. The automorphism group ${\rm Aut}(X_{F})$ is finite and equal to its subgroup ${\rm Lin}(X_{F})$ of linear automorphisms if $(n,d)\neq(1,3),(2,4)$ (see [MM63] for $n\geq 2$ and [Cha78] for $n=1$). See Section 2 for some notations on groups. We use ${\rm Aut}(F)$ to denote the group of linear transformations that preserve $F$. Then clearly $\mu_{n+2,d}$ is a normal subgroup of ${\rm Aut}(F)$ and the following exact sequence gives a basic relation between ${\rm Lin}(X_{F})$ and ${\rm Aut}(F)$:
$1\rightarrow\mu_{n+2,d}\longrightarrow{\rm Aut}(F)\longrightarrow{\rm Lin}(X_{F})\rightarrow 1$ (1.1)
@@
Bounding the size of ${\rm Aut}(X_{F})$, ${\rm Lin}(X_{F})$ or ${\rm Aut}(F)$ is almost the same thing and has a long history (see [OS78, Section 6] for some historical remarks). Bott–Tate and Orlik–Solomon showed that there exists an upper bound for the order $|{\rm Aut}(F)|$ depending only on $n$ and $d$ (see [OS78, Corollary (2.7)]). Finding an effective bound in terms of $n$ and $d$ is important. We use $X_{d}^{n}$ to denote the Fermat hypersurface in ${\mathbb{P}}^{n+1}$ of degree $d$ defined by Fermat polynomial $F_{d}^{n}:=x_{1}^{d}+x_{2}^{d}+\cdots+x_{n+2}^{d}$. It is known that ${\rm Aut}(X_{d}^{n})$ is isomorphic to $C_{d}^{n+1}\rtimes S_{n+2}$ and $|{\rm Aut}(X_{d}^{n})|=d^{n+1}\cdot(n+2)!$ if $(n,d)\neq(1,3),(2,4)$ (see [HS80], [Shi88]). By a classical theorem of Hurwitz, the automorphism group of a complex curve of genus $g\geq 2$ is of order at most $84(g-1)$. This implies that for a plane curve $X_{F}$ with $d\geq 4$, $|{\rm Aut}(X_{F})|\leq 42(d-3)d$. Automorphism groups of plane curves have been extensively studied by many people. For $n=1$ and $d\geq 4$, plane curves $X_{F}$ with maximum $|{\rm Aut}(X_{F})|$ are isomorphic to Fermat curves $X_{d}^{1}$ (resp. Klein quartic, resp. Wiman sextic) if $d\neq 4,6$ (resp. $d=4$, resp. $d=6$) (see [Pam13], [Har19] and references therein). For $n\geq 2$, Howard–Sommese [HS80, Theorem 2] proved that there is a constant $k_{n}$ depending only on $n$ such that $|{\rm Aut}(X_{F})|\leq k_{n}\cdot d^{n+1}$ if $(n,d)\neq(2,4)$. Xiao ([Xia94], [Xia95]) proved that a minimal smooth projective surface $S$ of general type has at most $42^{2}K_{S}^{2}$ automorphisms. Consequently, for a surface $X_{F}\subset{\mathbb{P}}^{3}$ with $d\geq 5$, $|{\rm Aut}(X_{F})|\leq 42^{2}(d-4)^{2}d$. For $2\leq n\leq 5$, Fermat cubic $n$-fold has the largest possible order for the automorphism group among all smooth cubic $n$-folds (see [Seg42], [Hos97], [Dol12] for $n=2$; [WY20] for $n=3$; [LZ22], [YYZ23] for $n=4$; [YYZ23] for $n=5$). Fermat quintic $3$-fold has the largest possible order for the automorphism group among all smooth quintic $3$-folds ([OY19]). For some partial results on abelian subgroups of automorphism groups of smooth hypersurfaces of arbitrary dimension, see [CS95], [GL11], [GL13], [Zhe22], [GLM23], etc. However, to the best of our knowledge, for $n\geq 2$ except $(n,d)=(2,3),(2,4),(3,3),(4,3),(5,3)$ and $(3,5)$, the maximum order $|{\rm Aut}(X_{F})|$ is still unknown (see e.g. [Huy23, Remark 3.20], [GLMV24, Remark 2.11], [Ess24, Question 3.9]) and the classification of groups of linear automorphisms of hypersurfaces is in a rudimentary state as also mentioned in [Dol19, Page 6]. Our main result is to completely classify smooth hypersurfaces with automorphism groups of maximum order, which in particular gives an optimal upper bound for $|{\rm Aut}(X_{F})|$.
@@
The sextic $F$ and its automorphism group in the case $(n,d)=(4,6)$ were given by Todd [Tod50]. It seems that $X_{F}$ in the cases $(n,d)=(2,6),(2,12),(4,12)$ in the table above are previously unknown. It is known that automorphism groups of smooth hypersurfaces of general type are equal to their birational automorphism groups. Hacon–${\rm M^{c}Kernan}$–Xu [HMX13] showed that the number of birational automorphisms of a variety of general type $X$ is bounded by $c\cdot{\rm vol}(X,K_{X})$, where $c$ is a constant that only depends on the dimension of $X$, and they asked for finding an explicit bound for the constant $c$ ([HMX13, Question 1.2]). As a direct consequence of Theorem 1.1, we obtain the optimal values of such constants for smooth hypersurfaces of general type.
@@
Next we briefly explain ideas of the proof of Theorem 1.1. In fact, we completely classify smooth hypersurfaces $X_{F}$ with $|{\rm Aut}(X_{F})|\geq|{\rm Aut}(X_{d}^{n})|$ (Theorem 7.5), which clearly implies Theorem 1.1. In order to prove Theorem 7.5, by the exact sequence (1.1), it suffices to classify $X_{F}$ with $|{\rm Aut}(F)|\geq|{\rm Aut}(F_{d}^{n})|$. A famous theorem of Jordan says that, for every integer $r\geq 1$, there exists a constant depending only on $r$ such that any finite subgroup of the general linear group ${\rm GL}(r,{\mathbb{C}})$ has an abelian normal subgroup of index at most this constant. Collins [Col07] gave the optimal values ${\rm J}(r)$ of such constants for all $r\geq 1$. By Jordan’s Theorem, as a finite linear group in ${\rm GL}(n+2,{\mathbb{C}})$, ${\rm Aut}(F)$ has a normal abelian subgroup $N$ of index at most ${\rm J}(n+2)$, and $|{\rm Aut}(F)|$ is bounded by the product $|N|\cdot{\rm J}(n+2)$. Hence bounding the size of $N$ leads to an upper bound of $|{\rm Aut}(F)|$. The known bounds for $|{\rm Aut}(F)|$ or $|{\rm Aut}(X_{F})|$ mentioned previously are obtained in this way or a similar way (see [OS78], [HS80]). However, it is often that when the index $[{\rm Aut}(F):N]$ gets bigger, the size of $N$ gets smaller. Thus the product of ${\rm J}(n+2)$ and the largest possible value for $|N|$ rarely gives an optimal bound for $|{\rm Aut}(F)|$ (see Remark 7.6). To overcome this difficulty, we introduce two new notions, canonical bound (Definition 3.5) and Fermat-test ratio (Definition 4.5). Roughly speaking, canonical bounds integrate bounding abelian subgroups of ${\rm Aut}(F)$ with bounding certain primitive groups called primitive constituents (see $\bf{Set\textrm{-}up}$ 3.1) of ${\rm Aut}(F)$, which turns out to be quite effective in our study of bounding $|{\rm Aut}(F)|$. More precisely, based on complete reducibility of linear representations of finite groups and the classical notion of primitive (projective) linear groups (see Section 3), we introduce two associated exact sequences (3.1)-(3.2) for any finite linear group. Combining the associated exact sequences with our bound for abelian subgroups of ${\rm Aut}(F)$ (Theorem 3.4), we show that $|{\rm Aut}(F)|$ is at most the canonical bound $B({\rm Aut}(F))$ of ${\rm Aut}(F)$ (Lemma 3.6). From this, we immediately get an optimal bound for $|{\rm Aut}(F)|$ in the case ${\rm Aut}(F)$ consisting of semi-permutations (Theorem 3.9).
@@
To handle the case ${\rm Aut}(F)$ consisting of not only semi-permutations, we use both canonical bounds and Fermat-test ratios. As a consequence of Jordan’s theorem, for each integer $r\geq 1$, there are only finitely many finite primitive groups in ${\rm PGL}(r,{\mathbb{C}})$ (up to conjugation), and Collins [Col08a, Theorem A] determined the maximum order ${\rm JC}(r)$ among such groups using the well-known classification of finite simple groups. Based on the definition of Fermat-test ratio and the explicit values of ${\rm JC}(r)$, we show that Fermat-test ratios have many nice properties (see Section 4) which are crucial for our classification. In particular, using Fermat-test ratios, we quickly prove that Theorem 7.5 holds if $n\geq 26$ or $d\geq 18$ (Theorem 4.13). Sections 5, 6 and 7 are devoted to the proof of Theorem 7.5 for the remaining (finitely many) pairs $(n,d)$ satisfying $n\leq 25$ and $d\leq 17$. For such pairs of $(n,d)$, we classify $X_{F}$ with $|{\rm Aut}(F)|\geq|{\rm Aut}(F_{d}^{n})|$ in three steps:
@@
In these steps, our classification is based on known classifications of finite primitive linear groups in small degrees (see Section 6) and close relations among smoothness and shape of the defining polynomial $F$, canonical bounds, and Fermat-test ratios (see Section 5). Especially, we show that monomials in $F$ involving different primitive constituents of ${\rm Aut}(F)$ considerably reduce the order $|{\rm Aut}(F)|$ (Lemmas 5.4, 5.5, 5.6, 5.7), which plays a key role in steps (2) and (3).
We conclude the introduction by posing a question closely related to our work. Collins [Col08b] obtained the optimal values of modular analogues of Jordan constants for finite linear groups. Dolgachev–Duncan [DD19] classified all possible automorphism groups of smooth cubic surfaces over an algebraically closed field of arbitrary characteristic. It would be interesting to study an analogue of Theorem 1.1 in positive characteristic.
@@ Acknowledgements
We would like to thank Professors Keiji Oguiso and Zhiwei Zheng for valuable conversations. This work is partially supported by the National Natural Science Foundation of China (No. 12171351, No. 12071337, No. 11921001).
@@ 2. Notation and conventions
(2.1)  Let $W$ be a complex vector space of dimension $r\geq 1$. Let $W^{*}$ denote the dual vector space of $W$ and $S(W^{*})$ be the symmetric algebra of $W^{*}$. Each element in $S(W^{*})$ can be viewed as a complex-valued function on $W$. Then ${\rm GL}(W)$ acts from left (resp. right) on $W$ (resp. $S(W^{*})$) via $(f,w)\mapsto f(w)$ (resp. $(F,f)\mapsto f(F)$). Here $f(F)$ is given by $f(F)(w)=F(f(w))$ for any $w\in W$. Note that the action of ${\rm GL}(W)$ on $W$ induces an action of ${\rm GL}(W)$ on ${\mathbb{P}}(W)$ given by $(f,[w])\mapsto[f(w)]$. Two elements $F,F^{\prime}$ in $S(W^{*})$ are called isomorphic to each other if they lie in the same orbit under the action by ${\rm GL}(W)$ on $S(W^{*})$.
@@
$(A,(w_{1},\ldots,w_{r}))\mapsto(\sum_{i=1}^{r}a_{1i}w_{i},\ldots,\sum_{i=1}^{r}a_{ri}w_{i})\,\,\text{ and }\,\,(F,A)\mapsto A(F):=F(\sum_{i=1}^{r}a_{1i}x_{i},\ldots,\sum_{i=1}^{r}a_{ri}x_{i})$
respectively. We call $(e_{1},\ldots,e_{r})$ (resp. $(e_{1}^{*},\ldots,e_{r}^{*})$) the underlying basis of ${\mathbb{C}}^{r}\cong W$ (resp. ${{\mathbb{C}}^{r}}^{*}\cong W^{*}$). In particular, in this paper, we always assume that the underlying bases of $W$ and $W^{*}$ are dual to each other, and if we choose a new basis $(e_{i}^{\prime})$ as the underlying basis of $W$, it is understood that the underlying basis of $W^{*}$ is changed to the dual basis $({e_{i}^{\prime}}^{*})$.
(2.3)  We say the underlying basis $(e_{i})$ of $W$ is compatible with a decomposition $W=W_{1}\oplus\cdots\oplus W_{s}$ if $e_{n_{1}+\cdots+n_{i-1}+j}\in W_{i}$ for $1\leq i\leq s$, $1\leq j\leq n_{i}$, where $n_{i}={\rm dim}(W_{i})$. In this case, for any integer $d\geq 0$, the space of $d$-forms $S^{d}(W^{*})\subset S(W^{*})$ admits a decomposition
@@
(2.4)  Note that if $A_{1},A_{2}\in{\rm GL}(r,{\mathbb{C}})$, we have $(A_{1}A_{2})(F)=A_{2}(A_{1}(F))$. The automorphism group ${\rm Aut}(F)$ of a homogeneous polynomial $F\in{\mathbb{C}}[x_{1},\ldots,x_{r}]$ of degree $d\geq 1$ is given by
@@
We say $F$ is smooth if $(0,\ldots,0)$ is the only solution for $\frac{\partial F}{\partial x_{1}}=\cdots=\frac{\partial F}{\partial x_{r}}=0$. In particular, if $r\geq 3$, a smooth form $F$ defines a smooth hypersurface $X_{F}\subset{\mathbb{P}}^{r-1}$ of dimension $r-2$. We denote by $\pi:{\rm GL}(r,{\mathbb{C}})\rightarrow{\rm PGL}(r,{\mathbb{C}})$ the projection map. For a subgroup $H$ of ${\rm GL}(r,{\mathbb{C}})$ (resp. ${\rm PGL}(r,{\mathbb{C}})$), we say $H$ preserves $F$ if $H=G$ (resp. $H=\pi(G)$) for some subgroup $G\subset{\rm Aut}(F)$. In this case, we call $F$ an $H$-invariant form of degree $d$. We say a monomial $\mathfrak{m}$ of degree $d$ is in $F$ (or $\mathfrak{m}\in F$) if the coefficient of $\mathfrak{m}$ is not zero in the expression of $F$. As in [OY19], for a finite subgroup $H_{1}\subset{\rm PGL}(r,{\mathbb{C}})$, we say $H_{1}$ admits a lifting (resp. an $F$-lifting) if there exists $\widetilde{H}_{1}\subset{\rm GL}(r,{\mathbb{C}})$ (resp. $\widetilde{H}_{1}\subset{\rm Aut}(F)$ ) with $\widetilde{H}_{1}\cong\pi(\widetilde{H}_{1})=H_{1}$.
@@
$\xi_{k}$ the $k$-th primitive root $e^{\frac{2\pi i}{k}}$ of unity, where $k$ is a positive integer;
@@
$\mu_{r,d}$ the subgroup of ${\rm GL}(r,{\mathbb{C}})$ of order $d$ generated by the scalar matrix $\xi_{d}I_{r}$;

### v2

@@ 1. Introduction
The purpose of this paper is to study smooth hypersurfaces in complex projective spaces with large automorphism groups. Let $F\in{\mathbb{C}}[x_{1},\ldots,x_{n+2}]$ be a smooth homogeneous polynomial of degree $d\geq 3$ and let $X_{F}\subset{\mathbb{P}}^{n+1}$ denote the smooth hypersurface defined by $F$, where $n\geq 1$. An automorphism $f\in{\rm Aut}(X_{F})$ of $X_{F}$ is called linear with respect to the embedding $X_{F}\hookrightarrow{\mathbb{P}}^{n+1}$ if $f$ extends to an automorphism of ${\mathbb{P}}^{n+1}$, i.e., $f$ is given by a linear change of homogeneous coordinates. The automorphism group ${\rm Aut}(X_{F})$ is finite and equal to its subgroup ${\rm Lin}(X_{F})$ of linear automorphisms if $(n,d)\neq(1,3),(2,4)$ (see [MM63, Theorems 1 and 2] for $n\geq 2$ and [Cha78, Theorem 2] for $n=1$). We use ${\rm Aut}(F)$ to denote the group of linear transformations that preserve $F$ and $\mathcal{I}_{n+2,d}$ to denote the subgroup of the general linear group ${\rm GL}(n+2,{\mathbb{C}})$ consisting of $d$ scalar matrices. Then clearly $\mathcal{I}_{n+2,d}$ is a normal subgroup of ${\rm Aut}(F)$ and the following exact sequence gives a basic relation between ${\rm Lin}(X_{F})$ and ${\rm Aut}(F)$:
$1\rightarrow\mathcal{I}_{n+2,d}\longrightarrow{\rm Aut}(F)\longrightarrow{\rm Lin}(X_{F})\rightarrow 1$ (1.1)
@@
Bounding the size of ${\rm Aut}(X_{F})$, ${\rm Lin}(X_{F})$ or ${\rm Aut}(F)$ is almost the same thing and has a long history (see [OS78, Section 6] for some historical remarks). Bott–Tate and Orlik–Solomon showed that there exists an upper bound for the order $|{\rm Aut}(F)|$ depending only on $n$ and $d$ (see [OS78, Corollary (2.7)]). Finding an effective bound in terms of $n$ and $d$ is important. We use $X_{d}^{n}$ to denote the Fermat hypersurface in ${\mathbb{P}}^{n+1}$ of degree $d$ defined by Fermat polynomial $F_{d}^{n}:=x_{1}^{d}+x_{2}^{d}+\cdots+x_{n+2}^{d}$. It is known that ${\rm Aut}(X_{d}^{n})$ is isomorphic to a semidirect product $C_{d}^{n+1}\rtimes S_{n+2}$ and $|{\rm Aut}(X_{d}^{n})|=d^{n+1}\cdot(n+2)!$ if $(n,d)\neq(1,3),(2,4)$ (see [HS80, Page 147], [Shi88]). By a classical theorem of Hurwitz, the automorphism group of a complex curve of genus $g\geq 2$ is of order at most $84(g-1)$. This implies that for a plane curve $X_{F}$ with $d\geq 4$, $|{\rm Aut}(X_{F})|\leq 42(d-3)d$. Automorphism groups of plane curves have been extensively studied by many people. For $n=1$ and $d\geq 4$, plane curves $X_{F}$ with maximum $|{\rm Aut}(X_{F})|$ are isomorphic to Fermat curves $X_{d}^{1}$ (resp. Klein quartic, resp. Wiman sextic) if $d\neq 4,6$ (resp. $d=4$, resp. $d=6$) (see [Pam13, Theorem 1], [Har19, Theorem 2.5] and references therein). Andreotti [And50] was the first to establish explicit bounds for the order of the automorphism group of varieties of general type in higher dimensions. For $n\geq 2$, Howard–Sommese [HS80, Theorem 2] proved that there is a constant $k_{n}$ depending only on $n$ such that $|{\rm Aut}(X_{F})|\leq k_{n}\cdot d^{n+1}$ if $(n,d)\neq(2,4)$. Xiao ([Xia94, Theorem 1], [Xia95, Theorem 2]) proved that a minimal smooth projective surface $S$ of general type has at most $42^{2}K_{S}^{2}$ automorphisms. Consequently, for a surface $X_{F}\subset{\mathbb{P}}^{3}$ with $d\geq 5$, $|{\rm Aut}(X_{F})|\leq 42^{2}(d-4)^{2}d$. For $2\leq n\leq 5$, Fermat cubic $n$-fold has the largest possible order for the automorphism group among all smooth cubic $n$-folds (see [Seg42, §100], [Hos97, Theorem 5.3], [Dol12, Theorem 9.5.8] for $n=2$; [WY20, Theorem 1.1] for $n=3$; [LZ22, Corollary 6.14], [YYZ23, Theorem 1.2] for $n=4$; [YYZ23, Theorem 1.1] for $n=5$). Fermat quintic $3$-fold has the largest possible order for the automorphism group among all smooth quintic $3$-folds ([OY19, Theorem 2.2]). For some partial results on abelian subgroups of automorphism groups of smooth hypersurfaces of arbitrary dimension, see for instance [CS95, Theorem 0.1], [Sza96, Bézout Lemma], [GL11, Theorem 2.6], [GL13], [Zhe22, Theorem 4.8], [GLM23]. However, to the best of our knowledge, for $n\geq 2$ except $(n,d)=(2,3),(2,4),(3,3),(4,3),(5,3)$ and $(3,5)$, the maximum order $|{\rm Aut}(X_{F})|$ is still unknown (see e.g. [Huy23, Chapter 1, Remark 3.20], [GLMV24, Remark 2.11], [Ess24, Question 3.9]) and the classification of groups of linear automorphisms of hypersurfaces is in a rudimentary state as also mentioned in [Dol19, Page 6]. Our main result is to completely classify smooth hypersurfaces with automorphism groups of maximum order, which in particular gives an optimal upper bound for $|{\rm Aut}(X_{F})|$.
@@
The sextic $F$ and its automorphism group in the case $(n,d)=(4,6)$ were given by Todd [Tod50, Section 6]. It seems that $X_{F}$ in the cases $(n,d)=(2,6),(2,12),(4,12)$ in the table above are previously unknown. It is known that automorphism groups of smooth hypersurfaces of general type are equal to their birational automorphism groups. Hacon–${\rm M^{c}Kernan}$–Xu [HMX13, Theorem 1.1] showed that the number of birational automorphisms of a variety of general type $X$ is bounded from above by $c\cdot{\rm vol}(X,K_{X})$, where $c$ is a constant that only depends on the dimension of $X$, and they asked for finding an explicit bound for the constant $c$ ([HMX13, Question 1.2]). As a direct consequence of Theorem 1.1, we obtain the optimal values of such constants for smooth hypersurfaces of general type.
@@
Next we briefly explain the ideas of the proof of Theorem 1.1. In fact, we completely classify smooth hypersurfaces $X_{F}$ with $|{\rm Aut}(X_{F})|\geq|{\rm Aut}(X_{d}^{n})|$ (Theorem 7.5), which clearly implies Theorem 1.1. In order to prove Theorem 7.5, by the exact sequence (1.1), it suffices to classify $X_{F}$ with $|{\rm Aut}(F)|\geq|{\rm Aut}(F_{d}^{n})|$. A famous theorem of Jordan [Jor78] says that, for every integer $r\geq 1$, there exists a constant depending only on $r$ such that any finite subgroup of the general linear group ${\rm GL}(r,{\mathbb{C}})$ has an abelian normal subgroup of index at most this constant. Collins [Col07, Theorems A, B and D] gave the optimal values ${\rm J}(r)$ of such constants for all $r\geq 1$. By Jordan’s Theorem, as a finite linear group in ${\rm GL}(n+2,{\mathbb{C}})$, ${\rm Aut}(F)$ has a normal abelian subgroup $N$ of index at most ${\rm J}(n+2)$, and $|{\rm Aut}(F)|$ is bounded from above by the product $|N|\cdot{\rm J}(n+2)$. Hence bounding the size of $N$ leads to an upper bound of $|{\rm Aut}(F)|$. The known bounds for $|{\rm Aut}(F)|$ or $|{\rm Aut}(X_{F})|$ mentioned previously are obtained in this way or a similar way (see [OS78, Corollary 2.7], [HS80, §3]). However, it is often that when the index $[{\rm Aut}(F):N]$ gets bigger, the size of $N$ gets smaller. Thus the product of ${\rm J}(n+2)$ and the largest possible value for $|N|$ rarely gives an optimal bound for $|{\rm Aut}(F)|$ (see Remark 7.6). To overcome this difficulty, we introduce two new notions, canonical bound (Definition 3.6) and Fermat-test ratio (Definition 4.5). Roughly speaking, canonical bounds integrate bounding the size of abelian subgroups of ${\rm Aut}(F)$ with bounding the size of certain primitive groups called primitive constituents (see $\bf{Set\textrm{-}up}$ 3.2) of ${\rm Aut}(F)$, which turns out to be quite effective in our study of bounding $|{\rm Aut}(F)|$. More precisely, based on complete reducibility of linear representations of finite groups and the classical notion of primitive (projective) linear groups (see Section 3), we introduce two associated exact sequences (3.1)-(3.2) for any finite linear group. Combining the associated exact sequences with our bound for the order of abelian subgroups of ${\rm Aut}(F)$ (Theorem 3.5), we show that $|{\rm Aut}(F)|$ is at most the canonical bound $B({\rm Aut}(F))$ of ${\rm Aut}(F)$ (Lemma 3.7). We say an invertible matrix $A$ is semi-permutation if $A$ is a diagonal matrix up to permutation of columns. Using Theorem 3.5, we immediately get an optimal bound for $|{\rm Aut}(F)|$ in the case ${\rm Aut}(F)$ consisting of semi-permutations (Theorem 3.9).
@@
To handle the case ${\rm Aut}(F)$ consisting of not only semi-permutations, we use both canonical bounds and Fermat-test ratios. As a consequence of Jordan’s theorem, for each integer $r\geq 1$, there are only finitely many finite primitive groups in ${\rm PGL}(r,{\mathbb{C}})$ (up to conjugation), and Collins [Col08a, Theorem A] determined the maximum order ${\rm JC}(r)$ among such groups using the well-known classification of finite simple groups. Based on the definition of Fermat-test ratios and the explicit values of ${\rm JC}(r)$, we show that Fermat-test ratios have many nice properties (see Section 4) which are crucial for our classification. In particular, using Fermat-test ratios, we quickly prove that Theorem 7.5 holds if $n\geq 26$ or $d\geq 18$ (Theorem 4.14). Sections 5, 6 and 7 are devoted to the proof of Theorem 7.5 for the remaining (finitely many) pairs $(n,d)$ satisfying $n\leq 25$ and $d\leq 17$. For such pairs of $(n,d)$, we classify $X_{F}$ with $|{\rm Aut}(F)|\geq|{\rm Aut}(F_{d}^{n})|$ in three steps:
@@
In these steps, our classification is based on known classifications of finite primitive linear groups in small degrees (see Section 6) and close relations among smoothness and shape of the defining polynomial $F$, canonical bounds, and Fermat-test ratios (see Section 5). Especially, we show that the existence of monomials in $F$ closely related to different primitive constituents of ${\rm Aut}(F)$ considerably reduce the order $|{\rm Aut}(F)|$ (Lemmas 5.4, 5.5, 5.6, 5.7), which plays a key role in steps (2) and (3).
After our paper appeared on arXiv, Louis Esser and Jennifer Li informed us that they are now obtaining a similar result to our Theorem 1.1 but in a different method [EL24]. Both results rely on Collins’ work on bounding the size of finite primitive complex linear groups, though the approaches employed are otherwise different.
We conclude the introduction by posing a question closely related to our work. Collins [Col08b, Theorems A and A] obtained the optimal values of modular analogues of Jordan constants for finite linear groups. Dolgachev–Duncan [DD19, Theorem 1.1] classified all possible automorphism groups of smooth cubic surfaces over an algebraically closed field of arbitrary characteristic. It would be interesting to study an analogue of Theorem 1.1 in positive characteristic.
@@ Acknowledgements
We would like to thank Professors Keiji Oguiso and Zhiwei Zheng for valuable conversations. We would also like to thank the referees for helpful comments and suggestions. This work is partially supported by the National Natural Science Foundation of China (No. 12171351, No. 12071337, No. 11921001).
@@ 2. Notation and conventions
(2.1)  Let $W$ be a complex vector space of dimension $r\geq 1$. Let $W^{*}$ denote the dual vector space of $W$ and $S(W^{*})$ be the symmetric algebra of $W^{*}$. Each element in $S(W^{*})$ can be viewed as a complex-valued function on $W$. Then ${\rm GL}(W)$ acts from the left (resp. right) on $W$ (resp. $S(W^{*})$) via $(f,w)\mapsto f(w)$ (resp. $(F,f)\mapsto f(F)$). Here $f(F)$ is given by $f(F)(w)=F(f(w))$ for any $w\in W$. Note that the action of ${\rm GL}(W)$ on $W$ induces an action of ${\rm GL}(W)$ on ${\mathbb{P}}(W)$ given by $(f,[w])\mapsto[f(w)]$. Two elements $F,F^{\prime}$ in $S(W^{*})$ are called isomorphic to each other if they lie in the same orbit under the action of ${\rm GL}(W)$ on $S(W^{*})$.
@@
$(A,(w_{1},\ldots,w_{r}))\mapsto\big{(}\sum_{i=1}^{r}a_{1i}w_{i},\ldots,\sum_{i=1}^{r}a_{ri}w_{i}\big{)}\,\text{ and }\,(F,A)\mapsto A(F):=F\big{(}\sum_{i=1}^{r}a_{1i}x_{i},\ldots,\sum_{i=1}^{r}a_{ri}x_{i}\big{)}$
respectively. We call $(e_{1},\ldots,e_{r})$ (resp. $(e_{1}^{*},\ldots,e_{r}^{*})$) the underlying basis of ${\mathbb{C}}^{r}\cong W$ (resp. ${{\mathbb{C}}^{r}}^{*}\cong W^{*}$). This means that we identify the $e_{i}$’s with the standard basis of ${\mathbb{C}}^{r}$. In particular, in this paper, we always assume that the underlying bases of $W$ and $W^{*}$ are dual to each other, and if we choose a new basis $(e_{i}^{\prime})$ as the underlying basis of $W$, it is understood that the underlying basis of $W^{*}$ is changed to the dual basis $({e_{i}^{\prime}}^{*})$.
(2.3)  We say the underlying basis $(e_{i})$ of $W$ is compatible with a decomposition $W=W_{1}\oplus\cdots\oplus W_{s}$ if $e_{n_{1}+\cdots+n_{i-1}+j}\in W_{i}$ for $1\leq i\leq s$, $1\leq j\leq n_{i}$, where $n_{i}={\rm dim}(W_{i})$ and $n_{0}=0$. In this case, for any integer $d\geq 0$, the space of $d$-forms $S^{d}(W^{*})\subset S(W^{*})$ admits a decomposition
@@
(2.4)  Note that if $A_{1},A_{2}\in{\rm GL}(r,{\mathbb{C}})$, we have $(A_{1}A_{2})(F)=A_{2}(A_{1}(F))$. We often use the notions of forms and homogeneous polynomials interchangeably. The automorphism group ${\rm Aut}(F)$ of a homogeneous polynomial $F\in{\mathbb{C}}[x_{1},\ldots,x_{r}]$ of degree $d\geq 1$ is given by
@@
We say $F$ is smooth if $(0,\ldots,0)$ is the only solution for $\frac{\partial F}{\partial x_{1}}=\cdots=\frac{\partial F}{\partial x_{r}}=0$. In particular, if $r\geq 3$, a smooth form $F$ defines a smooth hypersurface $X_{F}\subset{\mathbb{P}}^{r-1}$ of dimension $r-2$. We denote by $\pi:{\rm GL}(r,{\mathbb{C}})\rightarrow{\rm PGL}(r,{\mathbb{C}})$ the projection map. For a subgroup $H$ of ${\rm GL}(r,{\mathbb{C}})$ (resp. ${\rm PGL}(r,{\mathbb{C}})$), we say $H$ preserves $F$ if $H=G$ (resp. $H=\pi(G)$) for some subgroup $G\subset{\rm Aut}(F)$. In this case, we call $F$ an $H$-invariant form of degree $d$. We say a monomial $\mathfrak{m}$ of degree $d$ is in $F$ (or $\mathfrak{m}\in F$) if the coefficient of $\mathfrak{m}$ is not zero in the expression of $F$. As in [OY19, Definition 4.4], for a finite subgroup $H_{1}\subset{\rm PGL}(r,{\mathbb{C}})$, we say $H_{1}$ admits a lifting (resp. an $F$-lifting) if there exists $\widetilde{H}_{1}\subset{\rm GL}(r,{\mathbb{C}})$ (resp. $\widetilde{H}_{1}\subset{\rm Aut}(F)$) such that $\widetilde{H}_{1}\cong\pi(\widetilde{H}_{1})$ and $\pi(\widetilde{H}_{1})=H_{1}$.
@@
$\xi_{k}$ the primitive $k$-th root $e^{\frac{2\pi i}{k}}$ of unity, where $k$ is a positive integer;
@@
$\mathcal{I}_{r,d}$ the subgroup of ${\rm GL}(r,{\mathbb{C}})$ of order $d$ generated by the scalar matrix $\xi_{d}I_{r}$;